# Physical-regularized Hierarchical Generative Model for Metallic Glass Structural Generation and Energy Prediction

### v1

@@ 3.4 Theoretical Contribution of Physics-Informed Regularizer
$\mathcal{E}_{\mathrm{train}}(\theta):=\sum_{i=1}^{N}w_{i}\bigl{(}\mathcal{A}\bm{u}(\mathbf{x}_{i};\theta)-v(\mathbf{x}_{i})\bigr{)}^{2},$
and let $\theta^{\star}$ be a minimizer. The latent test error is $\mathcal{E}_{\mathrm{test}}:=\|\bm{u}(\cdot;\theta)-\bm{u}(\cdot;\theta^{\star})\|_{\mathcal{U}}$.
Theorem 2 (Generalization bound).
Under Assumptions 12, we have
@@
Because both our energy- and RDF-based penalties satisfy the Lipschitz-stability condition (Assumption 1) and are evaluated in low-dimensional space ($D$ = 3), each physics-informed term directly drives down $\mathcal{E}_{\mathrm{train}}$. By minimizing these terms we therefore provably tighten the upper bound on $\mathcal{E}_{\mathrm{test}}$, improving generalization to unseen configurations. In this way, our physical regularizer are more than heuristics—they provide a principled mechanism for guaranteeing better predictive performance. This theoretical picture also has evidence from our experiments. Figure 3(c) visualizes the latent space of the full GlassVAE, whereas Fig. S2 shows the same model without energy and RDF regularizes. The baseline VAE collapses into a tight cluster that fails to separate high- and low-energy structures. By contrast, the physics-aware GlassVAE learns a more expressive, well-spread latent landscape that clearly distinguishes configurations by energy—exactly the discriminative structure predicted by the theorem.
4 Experiments and Results
To evaluate the performance of GlassVAE, we trained the model on molecular-dynamics (MD) trajectories of a Cu50Zr50 metallic glass and then assessed it on an unseen test set. The system comprised 108 atoms with a glass-transition temperature $T_{g}=700\,$K, with atomic interactions described using the embedded atom method (EAM) potential[59]. The system was first equilibrated in the liquid state at 2000 K for 2 ns, then quenched to 1000 K in 50 K intervals at a cooling rate of 100 K per 6 ns and further cooled down in 20 K intervals at the rate of 100 K per 60 ns, all using the NPT ensemble. The simulations consisted of 108 atoms, and an MD time step of 1 fs is used. Periodic boundary conditions are employed in all three directions. We employed trajectories under 700 K, 760 K, 840 K, 920 K, 1000 K, with a total samples number of 50,000. The in-detailed data pre-processing pipeline are introduced in Sec.A.5. In order to show the scalability of performance of GlassVAE, we also evaluated GlassVAE on two independent Cu50Zr50 metallic-glass trajectories at 720K and 960K, both initiated from distinct starting configurations (see details in Sec.4.4 and S3).
In summary, GlassVAE’s hierarchical latent space combined with physics-informed regularization provides the flexibility needed for complex glassy systems and outperforms standard VAE variants (see Table.1), achieving superior accuracy in both property prediction and graph reconstruction. We show the detailed experimental results below and discuss them in more detail in the next section.
@@ 4 Experiments and Results
Distance VAE (distance matrix as input) 11.85 0.89 11.85 0.77
@@
Single Graph-VAE (single latent space) 29.32 < 0 0.18 0.94 < 0.1
GraphVAE (no RDF) 0.37 > 0.99 0.14 0.93 < 0.1
GlassVAE (best, + RDF) 0.32 > 0.99 0.025 > 0.99 0.091
4.1 Graph Reconstruction
@@
4.2 Energy Prediction and Latent Space Exploration
Another core aim of GlassVAE is accurate potential-energy prediction. On the test set (Figs. 3a–b), our model achieves an $R^{2}>0.99$, producing energy estimates whose overall distribution is virtually indistinguishable from the ground truth; even within defined subgroups (Fig. S1), it maintains $R^{2}\approx 0.85$. Overall, we achieved a $RMSE=0.32$ among test dataset, which means on average a $\sim$ 0.025 eV reconstruction error on average. A UMAP projection of the latent codes (Fig. 3c), colored by true energy, reveals two well-separated clusters corresponding to configurations below and above the glass-transition temperature. In contrast, the VAE without physics-informed regularization (Fig. S2) collapses into a single undifferentiated cluster. These results demonstrate that jointly training on energy not only delivers state-of-the-art accuracy but also embeds meaningful energetic structure into the latent space, enabling clear separation of physical regimes in accordance with our theoretical findings (Sec. 3.4).
4.3 Sampling from the Latent Space
A key advantage of VAEs is their ability to generate novel samples. Here, we demonstrate these generative capability via both conditional and random sampling:
@@ 4.3 Sampling from the Latent Space
Random Sampling  Random generation involves sampling from a simple standard normal distribution and then transforming this sample using learned mean and variance parameters, effectively drawing from the complex distribution learned by the model. Specifically, we sample directly from the standard normal distribution (the prior). We can also introduce controlled variations by adding a small amount of scaled noise to this initial sample before decoding. This perturbed sample is then passed through the decoder network to produce a final structure that represents a smooth deviation from the unperturbed version. Fig. 4 (a - b) show the results of 100 randomly generated samples and their calculated potential energy distributions. Fig. 4 (a) visualizes the new structures in latent space (cluster near known trajectories), indicating the generation of similar but novel structures. Moreover, the calculated potential energies of these generated structures (Fig. 4 (b)) align well with the energy distribution of the training data, suggesting they are physically meaningful.
@@
Conditional Generation  Instead, we could also generate atomic configurations whose potential energies lie within a specified interval by performing gradient-based refinement directly in the GlassVAE latent space. Figure 5 illustrates this process for a target energy of $E=-4.87\pm 0.01\ \mathrm{eV/atom}$. The five generated structures are marked in latent space as red ${}^{\prime}\times^{\prime}$ symbols.
@@
Both conditional generation and latent space optimization leverage the co-training of the energy prediction head to structure the latent space. By enforcing an energy constraint, as we discussed earlier in Sec. 3.3, the generated configurations are more likely to be physically realistic and adhere to the desired potential energy characteristics. This approach provides flexibility in generation—either directly conditioning on a target energy or refining latent samples post hoc via gradient-based optimization.
4.4 Dimension Choice and Scalability
@@
The choice of latent dimension is critical (Fig. 6): energy-prediction and edge-reconstruction errors decrease sharply as the latent dimension increases after $dim=10$, before plateauing in the 16–32 range. A latent code in this latter range effectively captures most structural variability without superfluous model capacity. Furthermore, co-training the decoder with a differentiable energy head invests the latent manifold with thermodynamic significance. By using gradient-steering to guide latent codes into specific energy windows, GlassVAE can synthesize configurations of predetermined stability. This allows for the rapid exploration of low-energy basins, offering an alternative to computationally intensive molecular dynamics sampling.
Figure 6: Model performance vs Latent Dimension (a) Potential energy prediction RMSE vs Latent Dimension (b) Edge attribution (pairwise distances) RMSE vs Latent Dimension
@@
Although randomly picked snapshots from a single trajectory are structural different, testing on held-out data alone may not fully demonstrate GlassVAE’s scalability. To address this, we evaluated GlassVAE on two independent Cu50Zr50 metallic-glass trajectories at 720K and 960K, both initiated from distinct starting configurations. As shown in Fig. S3, the model’s energy predictions for these external trajectories still show accurate prediction patterns as original test set, confirming GlassVAE’s ability to generalize to simulations with markedly different structural evolution. We observe a modest accuracy drop on the new testing trajectories, which likely reflects the limited structural diversity of our training set—only five trajectories at distinct temperatures. Due to computational constraints, we restricted our experiments to these five temperature trajectories to demonstrate crystals or SMILES for molecules. To bridge this gap, we introduce GlassVAE’s capabilities. Incorporating additional trajectories and a wider range of conditions in future work should further enhance the model’s robustness.
5 Conclusion and Discussion
@@
Generative modeling of glassy materials has long been hindered by their disordered nature ant the lack of long-range periodicity, which precludes compact encodings such as CIF for crystals or SMILES for molecules. To bridge this gap, we introduce GlassVAE: a permutation-, translation-, and rotation-invariant graph variational autoencoder with hierarchical latent spaces. The multi-level embedding employed in this model is essential because a single latent space rarely offers sufficient flexibility to capture both the geometric complexity and property variability of glassy systems, making it challenging to achieve high-fidelity structure reconstruction and accurate energy prediction simultaneously (Table 1). We further bolster GlassVAE with physics-informed losses—energy regression, radial distribution function matching, and positional consistency—to ground its generative capabilities in physical reality.
@@
To our knowledge, GlassVAE is the first architecture for inorganic glasses to unify structure generation and property prediction within a single, thermodynamically meaningful latent space. In addition, we also demonstrate GlassVAE’s scalability to new Cu50Zr50 trajectories initialized from different configurations. While current benchmarks use limited data from a 108-atom system to showcase the initial promise, future work can extend to larger, more complex glasses, disentangle latent factors (e.g., composition, local strain), incorporate additional thermodynamic observables (free-energy proxies, elastic constants), and integrate active-learning loops with simulations to further enhance fidelity and guide sampling in high-dimensional configuration space. Overall, we believe our approach paves the way for accelerated discovery of disordered materials by enabling the simultaneous mapping of energy landscapes and the proposal of candidate structures for further computational or experimental investigations.
Acknowledgment
Appendix A Appendix / supplemental material
A.1 Supplemental Figures
@@
A.2 Algorithm
1
Input: Atomic configuration $(\mathbf{R},\mathbf{t},E)$ with $N$ atoms
Output: Latent $(\mu,\log\sigma^{2})$; reconstructions $\hat{\mathbf{R}},\hat{Y},\hat{\mathbf{e}}$; predicted $\hat{E}$
2
  1. 31.
@@ A.2 Algorithm
4    Init node embeddings $\rightarrow$ $L$ message-passing layers $\rightarrow$ global add-pool
@@
    • Positions $\;\hat{\mathbf{R}}=\mathbf{R}^{(0)}+f_{\Delta}(\mathbf{z})$ // MLP decoding $z$ into per-atom displacements
@@
    detail of loss terms are list in Sec. A.3
  • Algorithm 1 Physical-Regularized Graph VAE (concise)
    A.3 Training
@@
    • Training loop: for each batch,
@@
    Computational Resources and Experiment Settings
    All experiments and model training were run on the University of Wisconsin–Madison High Throughput Computing (HTC) cluster, using NVIDIA A100-SXM4-80GB GPUs. The compute nodes run LINUX with CUDA 11.4, Python 3.9, and PyTorch 2.0.1.
    A.4 Theorem 2 Full Demonstration
    $\mathcal{A}\bm{u}=v$ (7)
    , where $\mathcal{A}:\mathcal{U}\rightarrow\mathcal{V}$ is an arbitrary operator satisfying $||v||_{\mathcal{V}}<\infty$, and $||\mathcal{A}\bm{u}||_{\mathcal{V}}<\infty,\forall\,||\bm{u}||_{\mathcal{U}}<\infty$. As a concrete example of our setting, $\bm{u}(\cdot;\theta)$ is a neural network parametrized by $\theta$. $\mathcal{A}$ can be the operator that calculates fitted RDF by our neural networks, and $v$ encodes our priori in RDF from the physical law.
    Assumption 3.
    (Regularities of $\mathcal{A}$) For any $\bm{u},\bm{u}^{\prime}\in\mathcal{U}$,
    $||\bm{u}-\bm{u}^{\prime}||_{\mathcal{U}}\leq C^{(1)}||\mathcal{A}\bm{u}-\mathcal{A}\bm{u}^{\prime}||_{\mathcal{V}}$ (8)
    Assumption 4.
[…]
    $\bar{g}_{N}:=\sum_{i=1}^{N}w_{i}g(\bm{u}_{i})$ (9)
    , where $\bm{u}_{i}$ is the quadrature point with its weights $w_{i}$. For low-dimensional space, it is safe to assume that
@@ A.4 Theorem 2 Full Demonstration
    $\approx\sum_{i=1}^{N}w_{i}(\mathcal{A}\bm{u}(\bm{x}_{i},\theta)-v)^{2}$
    , where $\{\bm{x}_{i}\}_{i=1}^{N}$ are the sample points in the training set $\mathcal{T}$, with $\{w_{i}\}_{i=1}^{N}$ the quadrature weights. We also define the test error
    $\mathcal{E}_{\text{test}}:=||\bm{u}(\cdot;\theta)-\bm{u}(\cdot;\theta^{*})||_{\mathcal{U}}$ (12)
@@
    Proof.
    $\mathcal{E}_{\text{test}}$ $=||\bm{u}(\cdot;\theta)-\bm{u}(\cdot;\theta^{*})||_{\mathcal{U}}\leq C^{(1)}||\mathcal{A}\bm{u}-\mathcal{A}\bm{u}^{\prime}||_{\mathcal{V}}$ (14)
    $\leq C^{(1)}\left(\sum_{i=1}^{N}w_{i}(\mathcal{A}\bm{u}(\bm{x}_{i},\theta)-v)^{2}+C^{(2)}N^{-\alpha}\right)^{1/2}$
@@
    Corollary 1.
    Suppose the operator $\mathcal{A}$ satisfies (8), the quadrature error satisfies (10), and the neural networks are trained by minimizing (11). Furthermore, the validation points and test points are sampled from the same distribution as the training samples $\{\bm{x}_{i}\}_{i=1}^{N}$ in the training set. Then,
    ¯$\bar{\mathcal{E}}_{\text{test}}\leq C^{(1)}\left(\bar{\mathcal{E}}_{\text{train}}+\mathcal{E}_{\text{vg}}+\frac{1}{\sqrt{N}}\sqrt{\hat{\sigma}^{2}(\theta^{*})}\right),$ (15)
    where ¯$\bar{\mathcal{E}}_{\text{train}}$ and ¯$\bar{\mathcal{E}}_{\text{test}}$ is the average train error and test errors across training and test data sets during different training processes, defined as
@@
    , with $\int\cdot d\mathcal{T}$ and $\int\cdot d\mathcal{T}^{\prime\prime}$ represents aggregation across different training data sets $\mathcal{T}$ and test data sets $\mathcal{T}^{\prime\prime}$, respectively. ¯$\bar{\mathcal{E}}_{\text{valid}}$ represents the expected difference between the training samples and validation samples, defined as
    ¯$\bar{\mathcal{E}}_{\text{valid}}=\frac{1}{N}\int_{\mathbb{D}^{3N}}\sum_{j=1}|\mathcal{A}\bm{u}(\bm{x}_{1\cdots N}^{\mathcal{T}};\theta)-\mathcal{A}\bm{u}(\bm{x}_{1\cdots N}^{\mathcal{T}};\theta^{*})|d\mathcal{T}^{\prime}$ (17)
    $\hat{\sigma}^{2}(\theta^{*})$ $=\mathbb{E}\Biggl{(}\int_{\mathbb{R}^{3N}}|\mathcal{A}\bm{u}(\bm{x}_{1\cdots N}^{\mathcal{T}};\theta)-\mathcal{A}\bm{u}(\bm{x}_{1}^{N};\theta^{*})|\,d\bm{x}_{1\cdots N}^{\mathcal{T}}$ (18)
    $\quad\quad-\int_{\mathcal{T}}\int_{\mathbb{D}^{3N}}|\mathcal{A}\bm{u}(\bm{x}_{1\cdots N}^{\mathcal{T}};\theta)-\mathcal{A}\bm{u}(\bm{x}_{1}^{N};\theta^{*})|\,d\bm{x}_{1\cdots N}^{\mathcal{T}}\,d\mathcal{T}\Biggr{)}^{2}.$
    Proof.
    ¯$\bar{\mathcal{E}}_{\text{test}}$ $=\mathbb{E}||\bm{u}(\cdot;\theta)-\bm{u}(\cdot;\theta^{*})||_{\mathcal{U}}\leq C^{(1)}\mathbb{E}||\mathcal{A}\bm{u}-\mathcal{A}\bm{u}^{\prime}||_{\mathcal{V}}$ (19)
    $\leq C^{(1)}\left(\mathbb{E}\left(\mathbb{E}(||\mathcal{A}\bm{u}-\mathcal{A}\bm{u}^{\prime}||_{\mathcal{V}})|\mathcal{T}\right)+\bar{\mathcal{E}}_{\text{valid}}-\bar{\mathcal{E}}_{\text{valid}}+\bar{\mathcal{E}}_{\text{test}}-\bar{\mathcal{E}}_{\text{test}}\right)$
    $\leq C^{(1)}\left(\bar{\mathcal{E}}_{\text{train}}+\mathcal{E}_{\text{vg}}+\frac{1}{\sqrt{N}}\sqrt{\mathbb{E}\left(\left|\mathbb{E}\left(||\mathcal{A}\bm{u}-\mathcal{A}\bm{u}^{\prime}||_{\mathcal{V}}|\mathcal{T}\right)-\bar{\mathcal{E}}_{V}\right|^{2}\right)}\right)$
    , and $\hat{\sigma}^{2}(\theta^{*})$ estimates $\mathbb{E}\left(\left|\mathbb{E}\left(||\mathcal{A}\bm{u}-\mathcal{A}\bm{u}^{\prime}||_{\mathcal{V}}|\mathcal{T}\right)-\bar{\mathcal{E}}_{V}\right|^{2}\right)$.
    A.5 Data Pre-processing
    Before model training, we convert raw simulation outputs into a graph-based dataset in four stages. First, we parse LAMMPS files to extract atomic coordinates and element labels, and we read corresponding potential energies from the EAM output and get $\{\mathbf{r}_{i},t_{i},E\}$ To ensure balanced sampling across thermodynamic states, we randomly select up to 5000 snapshots per temperature. Next, we normalize the energy values to a 0–100 scale, avoiding division by zero when the range is degenerate. Each configuration is then transformed into a PyTorch-Geometric Data object: atom types become one-hot node features, interatomic displacements (with periodic boundary corrections) and distances form edge attributes, and we store true positions for later radial-distribution-function regularization.

### v2

@@ Introduction
$\mathcal{E}_{\mathrm{train}}(\theta):=\sum_{i=1}^{N}w_{i}\bigl(\mathcal{A}\boldsymbol{u}(\mathbf{x}_{i};\theta)-v(\mathbf{x}_{i})\bigr)^{2},$
and let $\theta^{\star}$ be a minimizer. The latent test error is $\mathcal{E}_{\mathrm{test}}:=\|\boldsymbol{u}(\cdot;\theta)-\boldsymbol{u}(\cdot;\theta^{\star})\|_{\mathcal{U}}$.
Theorem 2 (Generalization bound)
Under Assumptions Theoretical Contribution of Physics-Informed RegularizerTheoretical Contribution of Physics-Informed Regularizer, we have
@@
Distance VAE (distance matrix as input) 11.85 0.89 0.24 0.77
@@
Single Graph-VAE (single latent space) 29.32 << 0 0.18 0.93 << 0.1
GraphVAE (no RDF) 0.37 >> 0.99 0.14 0.94 << 0.1
GlassVAE (best, + RDF) 0.32 >> 0.99 0.025 >> 0.99 0.091
Because both our energy‐ and RDF‐based penalties satisfy the Lipschitz‐stability condition (Assumption 1) and are evaluated in low-dimensional space ($D$ = 3), each physics‐informed term directly drives down $\mathcal{E}_{\mathrm{train}}$. By minimizing these terms we therefore provably tighten the upper bound on $\mathcal{E}_{\mathrm{test}}$, improving generalization to unseen configurations. In this way, our physical regularizer are more than heuristics—they provide a principled mechanism for guaranteeing better predictive performance. This theoretical picture also has evidence from our experiments. Figure 2(c) visualizes the latent space of the full GlassVAE, whereas Fig. S2 shows the same model without energy and RDF regularizes. The baseline VAE collapses into a tight cluster that fails to separate high- and low-energy structures. By contrast, the physics-aware GlassVAE learns a more expressive, well-spread latent landscape that clearly distinguishes configurations by energy—exactly the discriminative structure predicted by the theorem.
Experiments and Results
To evaluate the performance of GlassVAE, we trained the model on molecular‐dynamics (MD) trajectories of a Cu50Zr50 metallic glass and then assessed it on an unseen test set. The system comprised 108 atoms with a glass‐transition temperature $T_{g}=700\,$K, with atomic interactions described using the embedded atom method (EAM) potential. The system was first equilibrated in the liquid state at 2000 K for 2 ns, then quenched to 1000 K in 50 K intervals at a cooling rate of 100 K per 6 ns and further cooled down in 20 K intervals at the rate of 100 K per 60 ns, all using the NPT ensemble. The simulations consisted of 108 atoms, and an MD time step of 1 fs is used. Periodic boundary conditions are employed in all three directions. We employed trajectories under 700 K, 760 K, 840 K, 920 K, 1000 K, with a total samples number of 50,000. The in-detailed data pre-processing pipeline are introduced in Sec.Data Pre-processing. In order to show the scalability of performance of GlassVAE, we also evaluated GlassVAE on two independent Cu50Zr50 metallic‑glass trajectories at 720K and 960K, both initiated from distinct starting configurations (see details in Sec.Dimension Choice and Scalability Discussion and S3).
In summary, GlassVAE’s hierarchical latent space combined with physics‑informed regularization provides the flexibility needed for complex glassy systems and outperforms standard VAE variants (see Table.1), achieving superior accuracy in both property prediction and graph reconstruction. We show the detailed experimental results below and discuss them in more detail in the next section.
Energy Prediction and Latent Space Exploration
Another core aim of GlassVAE is accurate potential‐energy prediction. On the test set (Figs. 2a–b), our model achieves an $R^{2}>0.99$, producing energy estimates whose overall distribution is virtually indistinguishable from the ground truth; even within defined subgroups (Fig. S1), it maintains $R^{2}\approx 0.85$. Overall, we achieved a $RMSE=0.32$ among test dataset, which means on average a $\sim$ 0.025 eV reconstruction error on average. A UMAP projection of the latent codes (Fig. 2c), colored by true energy, reveals two well‐separated clusters corresponding to configurations below and above the glass‐transition temperature (mcinnes_umap_2020). In contrast, the VAE without physics‐informed regularization (Fig. S2) collapses into a single undifferentiated cluster. These results demonstrate that jointly training on energy not only delivers state‐of‐the‐art accuracy but also embeds meaningful energetic structure into the latent space, enabling clear separation of physical regimes in accordance with our theoretical findings (Sec. Theoretical Contribution of Physics-Informed Regularizer).
Graph Reconstruction
@@
Sampling from the Latent Space
A key advantage of VAEs is their ability to generate novel samples. Here, we demonstrate these generative capability via both conditional and random sampling: Random Sampling Random generation involves sampling from a simple standard normal distribution and then transforming this sample using learned mean and variance parameters, effectively drawing from the complex distribution learned by the model. Specifically, we sample directly from the standard normal distribution (the prior). We can also introduce controlled variations by adding a small amount of scaled noise to this initial sample before decoding. This perturbed sample is then passed through the decoder network to produce a final structure that represents a smooth deviation from the unperturbed version. Fig. 4 (a - b) show the results of 100 randomly generated samples and their calculated potential energy distributions. Fig. 4 (a) visualizes the new structures in latent space (cluster near known trajectories), indicating the generation of similar but novel structures. Moreover, the calculated potential energies of these generated structures (Fig. 4 (b)) align well with the energy distribution of the training data, suggesting they are physically meaningful.
@@
Conditional Generation Instead, we could also generate atomic configurations whose potential energies lie within a specified interval by performing gradient‐based refinement directly in the GlassVAE latent space. Figure 5 illustrates this process for a target energy of $E=-4.87\pm 0.01\ \mathrm{eV/atom}$. The five generated structures are marked in latent space as red $^{\prime}\times^{\prime}$ symbols.
@@
Both conditional generation and latent space optimization leverage the co-training of the energy prediction head to structure the latent space. By enforcing an energy constraint, as we discussed earlier in Sec. Generation, the generated configurations are more likely to be physically realistic and adhere to the desired potential energy characteristics. This approach provides flexibility in generation—either directly conditioning on a target energy or refining latent samples post hoc via gradient-based optimization.
Dimension Choice and Scalability Discussion
@@
The choice of latent dimension is critical (Fig. S4): energy-prediction and edge-reconstruction errors decrease sharply as the latent dimension increases after $dim=10$, before plateauing in the 16–32 range. A latent code in this latter range effectively captures most structural variability without superfluous model capacity. Furthermore, co-training the decoder with a differentiable energy head invests the latent manifold with thermodynamic significance. By using gradient-steering to guide latent codes into specific energy windows, GlassVAE can synthesize configurations of predetermined stability. This allows for the rapid exploration of low-energy basins, offering an alternative to computationally intensive molecular dynamics sampling.
@@
Although randomly picked snapshots from a single trajectory are structural different, testing on held‑out data alone may not fully demonstrate GlassVAE’s scalability. To address this, we evaluated GlassVAE on two independent Cu50Zr50 metallic‑glass trajectories at 720K and 960K, both initiated from distinct starting configurations. As shown in Fig. S3, the model’s energy predictions for these external trajectories still show accurate prediction patterns as original test set, confirming GlassVAE’s ability to generalize to simulations with markedly different structural evolution. We observe a modest accuracy drop on the new testing trajectories, which likely reflects the limited structural diversity of our training set—only five trajectories at distinct temperatures. Due to computational constraints, we restricted our experiments to these five temperature trajectories to demonstrate crystals or SMILES for molecules. To bridge this gap, we introduce GlassVAE’s capabilities. Incorporating additional trajectories and a wider range of conditions in future work should further enhance the model’s robustness.
Conclusion and Discussion
@@
Generative modeling of glassy materials has long been hindered by their disordered nature ant the lack of long‑range periodicity, which precludes compact encodings such as CIF for crystals or SMILES for molecules. To bridge this gap, we introduce GlassVAE: a permutation‑, translation‑, and rotation‑invariant graph variational autoencoder with hierarchical latent spaces. The multi‑level embedding employed in this model is essential because a single latent space rarely offers sufficient flexibility to capture both the geometric complexity and property variability of glassy systems, making it challenging to achieve high‐fidelity structure reconstruction and accurate energy prediction simultaneously (Table 1). We further bolster GlassVAE with physics‐informed losses—energy regression, radial distribution function matching, and positional consistency—to ground its generative capabilities in physical reality.
@@
To our knowledge, GlassVAE is the first architecture for inorganic glasses to unify structure generation and property prediction within a single, thermodynamically meaningful latent space. In addition, we also demonstrate GlassVAE’s scalability to new Cu50Zr50 trajectories initialized from different configurations. While current benchmarks use limited data from a 108‑atom system to showcase the initial promise, future work can extend to larger, more complex glasses, disentangle latent factors (e.g., composition, local strain), incorporate additional thermodynamic observables (free‑energy proxies, elastic constants), and integrate active‑learning loops with simulations to further enhance fidelity and guide sampling in high‑dimensional configuration space. Overall, we believe our approach paves the way for accelerated discovery of disordered materials by enabling the simultaneous mapping of energy landscapes and the proposal of candidate structures for further computational or experimental investigations.
Appendices
@@
Figure S4: Model performance vs Latent Dimension (a) Potential energy prediction RMSE vs Latent Dimension (b) Edge attribution (pairwise distances) RMSE vs Latent Dimension
Algorithm
Algorithm 1 Physical‑Regularized Graph VAE (concise)
Atomic configuration $(\mathbf{R},\mathbf{t},E)$ with $N$ atoms Latent $(\mu,\log\sigma^{2})$; reconstructions $\hat{\mathbf{R}},\hat{Y},\hat{\mathbf{e}}$; predicted $\hat{E}$
  1. 1.
@@
     Init node embeddings $\rightarrow$ $L$ message‑passing layers $\rightarrow$ global add‑pool
@@
    • Positions $\hat{\mathbf{R}}=\mathbf{R}^{(0)}+f_{\Delta}(\mathbf{z})$
@@
    detail of loss terms are list in Sec. Training
Training
@@
Computational Resources
All experiments and model training were run on the High Throughput Computing (HTC) cluster, using NVIDIA A100‑SXM4‑80GB GPUs. The compute nodes run LINUX with CUDA 11.4, Python 3.9, and PyTorch 2.0.1.
MD Experiment Settings
Theorem 2 Full Demonstration
$\mathcal{A}\boldsymbol{u}=v$ (7)
, where $\mathcal{A}:\mathcal{U}\rightarrow\mathcal{V}$ is an arbitrary operator satisfying $v||_{\mathcal{V}}<\infty$, and $\mathcal{A}\boldsymbol{u}||_{\mathcal{V}}<\infty,\forall\,||\boldsymbol{u}||_{\mathcal{U}}<\infty$. As a concrete example of our setting, $\boldsymbol{u}(\cdot;\theta)$ is a neural network parametrized by $\theta$. $\mathcal{A}$ can be the operator that calculates fitted RDF by our neural networks, and $v$ encodes our priori in RDF from the physical law.
(Assumption) Regularities of $\mathcal{A}$
For any $\boldsymbol{u},\boldsymbol{u}^{\prime}\in\mathcal{U}$,
@@
(Assumption) Quadrature error
For an arbitrary function $g\in\mathcal{V}$, we define its integral $\bar{g}:=\int g(\boldsymbol{u})d\boldsymbol{u}$, and the corresponding quadrature approximation
$\bar{g}_{N}:=\sum_{i=1}^{N}w_{i}g(\boldsymbol{u}_{i})$ (9)
, where $\boldsymbol{u}_{i}$ is the quadrature point with its weights $w_{i}$. For low-dimensional space, it is safe to assume that
@@
$\approx\sum_{i=1}^{N}w_{i}(\mathcal{A}\boldsymbol{u}(\boldsymbol{x}_{i},\theta)-v)^{2}$
, where $\{\boldsymbol{x}_{i}\}_{i=1}^{N}$ are the sample points in the training set $\mathcal{T}$, with $\{w_{i}\}_{i=1}^{N}$ the quadrature weights. We also define the test error
[…]
@@
[…]
$\leq C^{(1)}\left(\sum_{i=1}^{N}w_{i}(\mathcal{A}\boldsymbol{u}(\boldsymbol{x}_{i},\theta)-v)^{2}+C^{(2)}N^{-\alpha}\right)^{1/2}$
@@
Suppose the operator $\mathcal{A}$ satisfies (8), the quadrature error satisfies (10), and the neural networks are trained by minimizing (11). Furthermore, the validation points and test points are sampled from the same distribution as the training samples $\{\boldsymbol{x}_{i}\}_{i=1}^{N}$ in the training set. Then,
$\bar{\mathcal{E}}_{\text{test}}\leq C^{(1)}\left(\bar{\mathcal{E}}_{\text{train}}+\mathcal{E}_{\text{vg}}+\frac{1}{\sqrt{N}}\sqrt{\hat{\sigma}^{2}(\theta^{*})}\right),$ (15)
where $\bar{\mathcal{E}}_{\text{train}}$ and $\bar{\mathcal{E}}_{\text{test}}$ is the average train error and test errors across training and test data sets during different training processes, defined as
$\bar{\mathcal{E}}_{\text{train}}$ $=\int_{\mathbb{R}^{3N}}\mathcal{E}_{\text{train}}(\boldsymbol{x}_{1\cdots N}^{\mathcal{T}})d\mathcal{T},\quad\bar{\mathcal{E}}_{\text{test}}$ $=\int_{\mathbb{R}^{3N}}\mathcal{E}_{\text{test}}(\boldsymbol{x}_{1\cdots N}^{\mathcal{T}})d\mathcal{T}^{\prime\prime}$ (16)
, with $\int\cdot d\mathcal{T}$ and $\int\cdot d\mathcal{T}^{\prime\prime}$ represents aggregation across different training data sets $\mathcal{T}$ and test data sets $\mathcal{T}^{\prime\prime}$, respectively. $\bar{\mathcal{E}}_{\text{valid}}$ represents the expected difference between the training samples and validation samples, defined as
$\bar{\mathcal{E}}_{\text{valid}}=\frac{1}{N}\int_{\mathbb{D}^{3N}}\sum_{j=1}|\mathcal{A}\boldsymbol{u}(\boldsymbol{x}_{1\cdots N}^{\mathcal{T}};\theta)-\mathcal{A}\boldsymbol{u}(\boldsymbol{x}_{1\cdots N}^{\mathcal{T}};\theta^{*})|d\mathcal{T}^{\prime}$ (17)
$\hat{\sigma}^{2}(\theta^{*})$ $=\mathbb{E}\Biggl(\int_{\mathbb{R}^{3N}}|\mathcal{A}\boldsymbol{u}(\boldsymbol{x}_{1\cdots N}^{\mathcal{T}};\theta)-\mathcal{A}\boldsymbol{u}(\boldsymbol{x}_{1}^{N};\theta^{*})|\,d\boldsymbol{x}_{1\cdots N}^{\mathcal{T}}$ (18)
$\quad\quad-\int_{\mathcal{T}}\int_{\mathbb{D}^{3N}}|\mathcal{A}\boldsymbol{u}(\boldsymbol{x}_{1\cdots N}^{\mathcal{T}};\theta)-\mathcal{A}\boldsymbol{u}(\boldsymbol{x}_{1}^{N};\theta^{*})|\,d\boldsymbol{x}_{1\cdots N}^{\mathcal{T}}\,d\mathcal{T}\Biggr)^{2}.$
@@
[…]
$\leq C^{(1)}\left(\bar{\mathcal{E}}_{\text{train}}+\mathcal{E}_{\text{vg}}+\frac{1}{\sqrt{N}}\sqrt{\mathbb{E}\left(\left|\mathbb{E}\left(||\mathcal{A}\boldsymbol{u}-\mathcal{A}\boldsymbol{u}^{\prime}||_{\mathcal{V}}|\mathcal{T}\right)-\bar{\mathcal{E}}_{V}\right|^{2}\right)}\right)$
, and $\hat{\sigma}^{2}(\theta^{*})$ estimates $\mathbb{E}\left(\left|\mathbb{E}\left(||\mathcal{A}\boldsymbol{u}-\mathcal{A}\boldsymbol{u}^{\prime}||_{\mathcal{V}}|\mathcal{T}\right)-\bar{\mathcal{E}}_{V}\right|^{2}\right)$.
Data Pre-processing
Before model training, we convert raw simulation outputs into a graph‐based dataset in four stages. First, we parse LAMMPS files to extract atomic coordinates and element labels, and we read corresponding potential energies from the EAM output and get $\{\mathbf{r}_{i},t_{i},E\}$ . To ensure balanced sampling across thermodynamic states, we randomly select up to 5000 snapshots per temperature. Next, we normalize the energy values to a 0–100 scale, avoiding division by zero when the range is degenerate. Each configuration is then transformed into a PyTorch‐Geometric Data object: atom types become one‐hot node features, interatomic displacements (with periodic boundary corrections) and distances form edge attributes, and we store true positions for later radial‐distribution‐function regularization.